OpenStudy (anonymous):

Evaluate the expression when x=6 and y=-6. -y+8x

OpenStudy (anonymous):

Plug in 6 for x and -6 for y: -(-6)+8(6) =6 + 48 =54

OpenStudy (anonymous):

okay so 54 is the answer?

OpenStudy (anonymous):

@girlnotonfire

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

okay can u help with more?

OpenStudy (anonymous):

I can try!

OpenStudy (anonymous):

that is actually all the problems for now thanks for ur help :)

OpenStudy (anonymous):

Sure no problem :)